# Selective labeling with false discovery rate control

### v1

@@ Abstract
Obtaining high-quality labels for large datasets is expensive, requiring massive annotations from human experts. While AI models offer a cost-effective alternative by predicting labels, their label quality is compromised by the unavoidable labeling errors. Existing methods mitigate this issue through selective labeling, where AI labels a subset and human labels the remainder. However, these methods lack theoretical guarantees on the quality of AI-assigned labels, often resulting in unacceptably high labeling error within the AI-labeled subset. To address this, we introduce Conformal Labeling, a novel method to identify instances where AI predictions can be provably trusted. This is achieved by controlling the false discovery rate (FDR), the proportion of incorrect labels within the selected subset. In particular, we construct a conformal $p$-value for each test instance by comparing AI models’ predicted confidence to those of calibration instances mislabeled by AI models. Then, we select test instances whose $p$-values are below a data-dependent threshold, certifying AI models’ predictions as trustworthy. We provide theoretical guarantees that Conformal Labeling controls the FDR below the nominal level, ensuring that a predefined fraction of AI-assigned labels is correct on average. Extensive experiments demonstrate that our method achieves tight FDR control with high power across various tasks, including image and text labeling, and LLM QA.
@@ 1 Introduction
Large-scale, high-quality labeled data is crucial for the machine learning pipelines (Deng et al., 2009). While experts could provide high-quality labels for moderately sized datasets, the growing size of modern datasets has made this approach prohibitively expensive. AI models offer a cost-effective alternative by predicting labels, bypassing the need for human experts. However, AI models are prone to labeling error (Northcutt et al., 2021; Tan et al., 2024). For example, empirical evidence demonstrates that even state-of-the-art LLMs exhibit high labeling error when used for text annotation Baumann et al. (2025). The labeling error inherent to AI models significantly compromises their label quality, hindering the deployment of AI labeling for production. To balance the trade-off between labeling cost and error, selective labeling has been a promising solution (Geifman & El-Yaniv, 2017; Wang et al., 2023) by combining AI predictions with expert annotations.
Prior work on selective labeling primarily designed heuristic methods (Wang et al., 2021; Bernhardt et al., 2022; Wang et al., 2024a) that rely on AI models to label high-confidence instances while deferring the rest to human experts. However, these methods do not provide any theoretical guarantees on label quality. To address this, a recent work (Candès et al., 2025) proposes probably approximately correct (PAC) labeling, which guarantees that the overall labeling error is controlled with high probability. They achieve the guarantee by counterbalancing the error of AI labels with the zero error of expert labels. While the guarantee is theoretically appealing, the labeling error of AI models can be unacceptably high, even reaching 100%, while the overall labeling error is controlled. These limitations motivate us to investigate how to provably guarantee the quality of AI-assigned labels.
In this work, we propose Conformal Labeling, which formulates the labeling problem as multiple hypotheses testing. This method allows us to select a subset of the unlabeled dataset for AI labeling and ensure the quality of AI-assigned labels by providing a rigorous guarantee on the false discovery rate (FDR). In particular, given a labeled calibration dataset, we compute conformal $p$-values for test instances through a rank-based comparison of their predicted confidence against those of calibration instances where AI predictions are incorrect. This ensures that $p$-values for mislabeled instances stochastically dominate the uniform distribution on [0, 1], while $p$-values for correctly labeled instances are concentrated near zero, allowing selection procedures to distinguish between them easily. Then, we select test instances whose $p$-values are less than or equal to a data-driven threshold, which is delicately set with the calibration dataset to control FDR at the desired level. We further theoretically prove that Conformal Labeling achieves valid FDR control under mild assumptions, ensuring that the expected proportion of incorrect labels in the selected subset is below the desired level. While the label quality of existing methods depends heavily on model performance, Conformal Labeling guarantees label quality regardless of the underlying model’s performance by strictly controlling the FDR.
@@
To validate our method, we conduct extensive experiments on various labeling tasks, including image labeling (ImageNet (Deng et al., 2009), ImageNet-V2 (Recht et al., 2019)), text labeling (stance on global warming (Luo et al., 2021), and misinformation (Gabriel et al., 2022)), and LLM QA (MedMCQA (Pal et al., 2022), MMLU (Hendrycks et al., 2021), MMLU-Pro (Wang et al., 2024a)) tasks. The results demonstrate that Conformal Labeling achieves high power with controlled FDR, indicating that AI models can label a large proportion of data with high quality. For example, Conformal Labeling can label 58.67% of the ImageNet dataset with ResNet-34 (He et al., 2016), while keeping the FDR below 10%. In comparison, a naive approach of using AI-assigned labels for the entire dataset results in a labeling error of over 25%. Moreover, through comprehensive ablation studies, we validate that Conformal Labeling is robust to the size of calibration datasets, and more powerful models enable better selection results.
@@
  • We propose Conformal Labeling, a novel method for identifying instances where AI predictions could be provably trusted. Regardless of AI models’ performance, Conformal Labeling guarantees the quality of AI-assigned labels by strictly controlling the FDR.
  • We theoretically prove that Conformal Labeling provides a strict quality guarantee for AI-assigned labels: it achieves valid FDR control, ensuring the expected proportion of incorrect labels is below a user-specified level.
  • We empirically show that Conformal Labeling significantly reduces the labeling cost while tightly controlling the FDR, through extensive experiments conducted on image labeling, text labeling, and LLM QA tasks with various models.
@@ 2 Preliminaries
In this work, we study the problem of identifying instances where AI predictions can be provably trusted. Here, we give a formulation of multi-class classification as an example. Let $\mathcal{X}$ denote the feature space and $\mathcal{Y}=\{1,\ldots,K\}$ denote the label space. The test dataset $\mathcal{D}_{\mathrm{test}}=\{X_{j}\}_{j=1}^{m}$ consists of $m$ data instances, sampled i.i.d. from a data distribution $\mathcal{P}_{\mathcal{X}}$. We denote the unseen ground-truth labels of instance $X_{j}$ as $Y_{j}$. Besides, we consider a pre-trained AI model $f:\mathcal{X}\to\mathbb{R}^{|\mathcal{Y}|}$ used to generate labels for the test dataset. For a given test instance $X$, the AI model predicts the label with the largest estimated probability $\hat{Y}=\arg\max_{y\in\mathcal{Y}}f_{y}(X)$, where $f_{y}(X)$ denotes the estimated class probability for class $y\in\mathcal{Y}$.
Since AI models are typically prone to labeling errors, we aim to select a large subset from the test dataset $\mathcal{D}_{\mathrm{label}}$ to control the portion of incorrect labels. Formally, our goal is to identify the largest subset of indices $\mathcal{R}\subseteq\{1,\cdots,m\}$ that controls the false discovery rate (FDR), defined as below:
$\text{FDR}=\mathbb{E}\left[\frac{|\mathcal{R}\cap\mathcal{H}_{0}|}{\max(|\mathcal{R}|,1)}\right],$
where $H_{0}=\{j\in\{1,\cdots,m\}:Y_{j}\neq\hat{Y}_{j}\}$ is the set of indices with incorrect predictions. For notation shorthand, we denote $m]=\{1,\cdots,m\}$ in the following. The FDR metric measures the expected proportion of mislabeled samples within the AI-labeled subset, illustrating the quality of AI-assigned labels by explicitly bounding the fraction of incorrect labels.
In addition to FDR control, we also expect AI models to label as many test instances as possible, which corresponds to maximizing power:
$\text{Power}=\mathbb{E}\left[\frac{|\mathcal{R}\cap\mathcal{H}_{1}|}{\max(|\mathcal{H}_{1}|,1)}\right],$ (1)
@@
In this work, we assume access to a small labeled calibration dataset $\mathcal{D}_{\mathrm{cal}}=\{(X_{i},Y_{i})\}_{i=1}^{n}$. For convenience, we denote the test dataset as $\mathcal{D}_{\mathrm{test}}=\{(X_{j},Y_{j})\}_{j=n+1}^{n+m}$, where $Y_{j}$ is not observed. Since the labeling cost of a small dataset by human annotators is typically affordable, this assumption is practical in the real world and is also adopted in prior work (Candès et al., 2025). Besides, we assume that examples of the test and calibration datasets are both drawn i.i.d. from the joint distribution $\mathcal{P}_{\mathcal{X}\mathcal{Y}}$, a common setting in selective labeling (Jung et al., 2024; Candès et al., 2025).
@@
Ensuring high-quality labels while reducing annotation costs has motivated extensive research on selective labeling. Prior work on selective labeling primarily focused on heuristic methods. For example, some studies design collaborative annotation frameworks that combine expert labels with LLM labels to streamline the annotation process (Li et al., 2023; Kim et al., 2024). Others propose domain-specific methods, such as meta-learning strategies for medical image labeling (Vrabac et al., 2022), annotation frameworks for text data (Duan & Lalor, 2023), and human–AI collaborative systems for object detection (Zhang et al., 2025). Although these heuristic approaches effectively reduce annotation costs, they lack formal guarantees on the label quality, which can result in unreliable labels when AI models perform poorly.
Probably approximately correct (PAC) labeling (Candès et al., 2025) addresses this limitation by providing a theoretical guarantee: the overall labeling error across the dataset is controlled with high probability. At its core, PAC labeling strategically collects zero-error expert labels for instances where the AI model exhibits the highest uncertainty, while relying on potentially imperfect AI predictions for more certain instances. This strategic allocation ensures that the dataset’s overall labeling error is small, as the zero-error expert annotations effectively counterbalance the error introduced by the AI-assigned labels. However, because the guarantee only applies at the aggregate dataset level, the subset labeled by AI may exhibit high labeling error. This limitation underscores a critical gap: existing selective labeling methods cannot ensure the quality of AI-assigned labels, hindering their reliable deployment in real-world applications. To address this, we explore methods to provably guarantee the quality of AI-assigned labels in selective labeling.
@@ 3.1 Conformal Labeling
Our previous section shows that existing methods cannot guarantee the quality of AI-assigned labels. To address this limitation, we propose Conformal Labeling, which identifies instances where AI predictions can be provably trusted by controlling the FDR. Our approach is composed of three primary steps: quantifying uncertainty, constructing conformal $p$-values, and thresholding.
@@
Our approach builds on a key insight: we should select instances where the model exhibits high confidence in its predictions. To quantify the model confidence, we define an uncertainty score $\mathcal{S}:\mathcal{X}\to\mathbb{R}$, where a higher value indicates greater model uncertainty. We note that in the conformal inference framework, this score function is also known as the non-conformity score function. For example, we employ $\mathcal{S}(X)=1-\max_{y\in\mathcal{Y}}f_{y}(X)$ as our uncertainty score function (Hendrycks & Gimpel, 2016). This score is a valid measure of uncertainty, since prior works establish that misclassified samples generally receive lower probability scores (i.e., $\max_{y\in\mathcal{Y}}f_{y}(X)$) than correctly classified ones (Hendrycks & Gimpel, 2016; Tu et al., 2024).
@@
$H_{j}^{0}:Y_{n+j}\neq\hat{Y}_{n+j}\quad\text{v.s.}\quad H_{j}^{1}:Y_{n+j}=\hat{Y}_{n+j},\quad\forall j=1,\cdots,m$
Rejecting the null hypothesis $H_{j}^{0}$ indicates that $(X_{n+j},\hat{Y}_{n+j})$ should be included in the subset, as it is deemed to be classified correctly. To construct the selection subset, we employ conformal p-value, which builds upon conformal inference framework (Vovk et al., 1999; 2005). The underlying intuition is that: a test instance $X$ is likely a misclassification if its uncertainty score $\mathcal{S}(X)$ is generally larger than the scores of instances that are known to be misclassified. Leveraging this idea, conformal $p$-value is computed through a rank-based comparison of $\mathcal{S}(X)$ against uncertainty scores of misclassified instances.
@@
1:Mislabeled Calibration set $\mathcal{D}^{0}_{\mathrm{cal}}=\{(X_{i},Y_{i})\}_{i=1}^{n_{0}}$, test instances $\{X_{n+j}\}_{j=1}^{m}$, pre-trained classifier $f$, calibration set size $n=|\mathcal{D}_{\mathrm{cal}}|$, FDR target $\alpha\in(0,1)$.
2:# 1. Compute uncertainty scores
3:for $i=1,\dots,n+m$ do
4: Compute $\mathcal{S}_{i}:=1-\max_{y\in\mathcal{Y}}f_{y}(X_{i})$.
5:end for
6:# 2. Construct conformal p-values
7:for $j=1,\dots,m$ do
8: Construct $\hat{p}_{j}$ according to equation 2.
9:end for
10:# 3. Thresholding
11:Compute $j^{*}=\max\left\{j:\hat{p}_{(j)}\leq\frac{\alpha j(n+1)}{m(n_{0}+1)}\right\}$, where $\hat{p}_{(j)}$ is the $j$-th smallest p-value.
12:Output: $\mathcal{R}=\{j:\hat{p}_{j}\leq\hat{p}_{(j^{*})}\}$.
Formally, for the calibration dataset $\mathcal{D}_{cal}$, we identify the subset $\mathcal{D}_{cal}^{0}\subseteq\mathcal{D}_{cal}$ where instances are misclassified by the AI model. For simplicity, we denote $\mathcal{D}_{cal}^{0}=\{(X_{i},Y_{i})\}_{i=1}^{n_{0}}$, and thus $Y_{i}\neq\hat{Y}_{i}$ for $i=1,\cdots,n_{0}$. We compute the uncertainty scores for the entire dataset $\{(X_{i},Y_{i})\}_{i=1}^{n+m}$: $\mathcal{S}_{i}=1-\max_{y\in\mathcal{Y}}f_{y}(X_{i})$. Then, the conformal $p$-value for the instance $X_{n+j}$ is computed by
$\hat{p}_{j}=\frac{\sum_{i=1}^{n_{0}}\mathbf{1}\{\mathcal{S}_{i}<\mathcal{S}_{n+j}\}+\left(1+\sum_{i=1}^{n_{0}}\mathbf{1}\{\mathcal{S}_{i}=\mathcal{S}_{n+j}\}\right)\cdot U_{j}}{n_{0}+1},$ (2)
where $U_{j}\sim\mathrm{Uniform}[0,1]$ are i.i.d. uniform random variable to randomize over ties when $S_{n+j}$ equals some $S_{i}$, ensuring a continuous conformal $p$-value. Here, $\hat{p}_{j}$ quantifies how extreme the uncertainty of $X_{n+j}$ is compared to the scores of misclassified instances, with a small $\hat{p}_{j}$ providing strong statistical evidence for correct prediction.
Standard results from conformal inference establish the validity of conformal $p$-value in practice: if the instance $X_{n+j}$ is misclassified, then the conformal $p$-value stochastically dominates the uniform distribution on $[0,1]$ (Bates et al., 2023; Jin & Candès, 2023): $\mathbb{P}\{\hat{p}_{j}\leq\alpha\mid H_{j}^{0}\text{ is true}\}\leq\alpha.$ This property indicates that the conformal $p$-values for misclassified instances are biased to be high, which allows us to set a threshold to flag potential correct instances, while controlling the FDR.
Thresholding.
After obtaining conformal $p$-values for test instances, we apply a thresholding rule inspired by the Benjamini–Hochberg (BH) procedure (Benjamini & Hochberg, 1995) to select a maximal subset $\mathcal{R}$ for AI labeling while controlling the FDR at level $\alpha$. The key idea is that we gradually increase the acceptance threshold until including additional samples would risk exceeding the desired labeling error. In particular, let $p_{(1)}\leq\ldots\leq p_{(m)}$ denote the ordered statistics of the $p$-values; the rejection set of our selection procedure applied to the conformal $p$-values is $\mathcal{R}=\{j:\hat{p}_{j}\leq\hat{p}_{(j^{*})}\}$, where
$j^{*}=\max\left\{j:\hat{p}_{(j)}\leq\frac{\alpha j(n+1)}{m(n_{0}+1)}\right\},$
[…]
Theoretical results.
We now provide a theoretical guarantee for Conformal Labeling. In the following theorem, we establish that Algorithm 1 controls the FDR below the desired level $\alpha$. This theorem ensures that AI predictions can be provably trusted in the selected subset, as the expected proportion of incorrect labels is controlled below $\alpha$. The proof is provided in Appendix B.2.

### v2

# Model-agnostic Selective Labeling with Provable Statistical Guarantees

@@ Abstract
Obtaining high-quality labels for large datasets is expensive, requiring massive annotations from human experts. While AI models offer a cost-effective alternative by predicting labels, their predictions inevitably contain errors. Existing methods mitigate this issue through selective labeling, where AI labels a subset and experts label the rest. However, these methods lack rigorous theoretical guarantees on the quality of AI-assigned labels, potentially leading to substantial labeling errors. To address this, we propose Conformal Labeling, a novel and model-agnostic method to identify a subset of unlabeled data with strict false discovery rate (FDR) control. In particular, we leverage a small labeled calibration set to construct conformal $p$-values for test data and estimate the accuracy of AI models. Then, we determine a data-dependent threshold for $p$-value selection, which automatically adapts to the estimated accuracy—higher accuracy permits a larger threshold. We provide theoretical guarantees that Conformal Labeling controls the FDR below the nominal level, ensuring that a predefined fraction of AI-assigned labels is correct in expectation. Extensive experiments across a wide range of models and datasets demonstrate that our method can achieve the target labeling quality with high power in classification and open-ended generation tasks.
Selective Labeling, Uncertainty Quantification, Conformal Inference
@@ 1 Introduction
Large-scale, high-quality labeled data is crucial for the machine learning pipelines (Deng et al., 2009). While human experts can provide reliable annotations for moderately sized datasets, the rapid growth of modern datasets has rendered this approach prohibitively expensive. AI models offer a cost-effective alternative by predicting labels. However, AI models are prone to labeling error (Northcutt et al., 2021; Tan et al., 2024), with empirical evidence showing that state-of-the-art LLMs exhibit high error in text annotation tasks (Baumann et al., 2025). The prediction errors compromise label quality, hindering the deployment of AI labeling in real-world applications. To balance the trade-off between labeling cost and error, selective labeling has been a promising solution (Li et al., 2023; Wang et al., 2023) by combining AI predictions with expert annotations.
Prior work on selective labeling primarily designed heuristic methods (Wang et al., 2021; Bernhardt et al., 2022; Wang et al., 2024a). These methods typically rely on model confidence, assigning high-confidence instances to AI models while deferring the remaining instances to human experts. However, these methods provide no theoretical guarantee on label quality. When the AI model is inaccurate or the confidence measure is unreliable, a substantial number of incorrect AI-assigned labels will be introduced, leading to a low-quality labeled dataset. Explicit accuracy guarantees directly address this issue by quantifying and controlling the quality of AI-assigned labels, ensuring that the resulting labeled dataset has a bounded error level pre-defined by users. This motivates us to design a model-agnostic method with provable accuracy guarantees on the AI-assigned labels.
In this work, we propose Conformal Labeling, a novel method to identify a subset of unlabeled data with provable FDR control. In particular, we leverage a small labeled calibration set to construct conformal $p$-values for test data and estimate the accuracy of the AI model. Motivated by our analysis of the conservatism of the Benjamini–Hochberg (BH) procedure (Benjamini & Hochberg, 1995), we apply the BH procedure at an adjusted significance level derived from the estimated model accuracy, which enables effective FDR control while achieving high selection power. We provide theoretical guarantees that Conformal Labeling controls the FDR with distribution-free, finite-sample guarantees. While the error rate of AI labeling in current methods depends heavily on model performance, Conformal Labeling can control the desired labeling error in expectation regardless of the underlying model’s performance.
@@
We empirically validate our method through extensive experiments conducted on three labeling tasks, including image labeling (ImageNet (Deng et al., 2009), ImageNet-V2 (Recht et al., 2019)), LLM QA (MedMCQA (Pal et al., 2022), MMLU (Hendrycks et al., 2021), MMLU-Pro (Wang et al., 2024a)), and LLM open-ended generation (MMLU-Redux (Gema et al., 2025), MATH-500 (Lightman et al., 2023) MATH-L5 (Hendrycks et al., 2021), Zebra-Logic (Lin et al., 2025), and $\text{HumanEval}^{+}$ (Liu et al., 2023)) tasks. The results demonstrate that Conformal Labeling achieves high power with controlled FDR, indicating that AI models can label a large proportion of data with high quality. For example, Conformal Labeling can label 63.73% of Zebra-Logic with Qwen3-4B-Thinking-2507 (Yang et al., 2025), while controlling the FDR at 9.85%. In comparison, using AI-assigned labels for the entire dataset results in labeling errors of over 10%. Moreover, through comprehensive analyses, we validate that Conformal Labeling is robust to the size of calibration datasets, and the logits-based uncertainty score performs better than the verbalized uncertainty score.
@@
  • We propose Conformal Labeling, a novel method for identifying a subset where AI predictions could be provably trusted. Regardless of AI models’ performance, Conformal Labeling guarantees the quality of AI-assigned labels by strictly controlling the FDR.
  • We theoretically prove that Conformal Labeling provides a strict quality guarantee for AI-assigned labels: it achieves FDR control, ensuring the expected portion of incorrect labels is below a user-specified level.
  • We empirically show that Conformal Labeling can label a large portion of data with AI models while tightly controlling the FDR, through extensive experiments conducted on image labeling, LLM QA, and LLM open-ended generation tasks with various models.
@@ 2 Preliminaries
We study the problem of identifying a subset where AI predictions can be provably trusted. Here, we give a formulation of multi-class classification as an example, while our framework also applies to open-ended generation. Let $\mathcal{X}$ denote the feature space and $\mathcal{Y}=\{1,\ldots,K\}$ denote the label space. The test dataset $\mathcal{D}_{\mathrm{test}}=\{X_{j}\}_{j=1}^{m}$ consists of $m$ data instances, sampled i.i.d. from a data distribution $\mathcal{P}_{\mathcal{X}}$. We denote the unseen ground-truth labels of instance $X_{j}$ as $Y_{j}$. Besides, we consider a pre-trained AI model $f:\mathcal{X}\to\mathbb{R}^{|\mathcal{Y}|}$ used to predict labels for the test dataset. For a given test instance $X$, the AI model predicts the label with the largest estimated probability $\widehat{Y}=\arg\max_{y\in\mathcal{Y}}f_{y}(X)$, where $f_{y}(X)$ denotes the estimated class probability for class $y\in\mathcal{Y}$.
Since AI models unavoidably make errors when used for labeling, we aim to select a large subset from the test dataset $\mathcal{D}_{\mathrm{label}}$ to control the portion of incorrect labels. Formally, our goal is to identify the largest subset of indices $\mathcal{R}\subseteq\{1,\cdots,m\}$ that controls the FDR, defined as below:
$\mathrm{FDR}=\mathbb{E}\left[\frac{|\mathcal{R}\cap\mathcal{H}_{0}|}{\max(|\mathcal{R}|,1)}\right],$ (1)
where $\mathcal{H}_{0}=\{j\in\{1,\cdots,m\}:Y_{j}\neq\widehat{Y}_{j}\}$ is the set of indices with incorrect predictions, and the expectation is taken over the randomness of the data. For notation shorthand, we denote $m]=\{1,\cdots,m\}$ in the following. The FDR metric measures the expected proportion of mislabeled samples within the AI-labeled subset, illustrating the quality of AI-assigned labels by explicitly bounding the expected fraction of incorrect labels.  [^1] The realized FDR might exceed the target level for a single run.
In addition to FDR control, we also expect AI models to label as many test instances as possible correctly, which corresponds to maximizing power:
$\mathrm{Power}=\mathbb{E}\left[\frac{|\mathcal{R}\cap\mathcal{H}_{1}|}{\max(|\mathcal{H}_{1}|,1)}\right],$ (2)
@@
In this work, we assume access to a small labeled calibration dataset $\mathcal{D}_{\mathrm{cal}}=\{(X_{i},Y_{i})\}_{i=1}^{n}$. For convenience, we denote the test dataset as $\mathcal{D}_{\mathrm{test}}=\{(X_{j},Y_{j})\}_{j=n+1}^{n+m}$, where $Y_{j}$ is not observed. Since the labeling cost of a small dataset by human annotators is typically affordable, this assumption is practical in the real world and is also adopted in prior work (Candès et al., 2025). Besides, we assume that examples of the test and calibration datasets are both drawn i.i.d. from the joint distribution $\mathcal{P}_{\mathcal{X}\mathcal{Y}}$. This assumption is realistic in our setting because we sample the calibration set from a large unlabeled dataset and leave the remaining data for testing. The process ensures that calibration data and test data are naturally i.i.d. In Appendix I.1, we test our proposed method against distribution shift to show that our proposed method is empirically robust to distribution shift.
@@
Ensuring high-quality labels while reducing annotation costs has motivated extensive research on selective labeling. Prior work on selective labeling primarily focused on heuristic methods. For example, several works explore human-LLM collaboration for annotation, including uncertainty-guided work allocation (Li et al., 2023), verifier-based quality assessment that routes low-confidence samples to humans (Wang et al., 2024b), and unsupervised collaboration between LLMs and small models (Xiao et al., 2023).
These methods balance labeling cost and quality by empirically selecting confidence thresholds to route samples between AI models and human annotators, yet provide no theoretical guarantees on the reliability of the AI-assigned labels. These methods implicitly assume well-calibrated confidence scores and sufficiently strong underlying model performance. However, commonly used confidence measures—including softmax probabilities from deep neural networks, logit-based confidence scores from LLMs, and LLM-generated verbalized confidence—are known to be poorly calibrated and overconfident (Guo et al., 2017; Chhikara, 2025), making them unreliable. Moreover, when the underlying model exhibits limited accuracy, even predictions assigned high confidence may be incorrect. Consequently, when the models are inaccurate or the confidence measure is unreliable, these methods will result in substantial labeling errors. These limitations motivate us to explore model-agnostic methods that can provably guarantee the quality of AI-assigned labels in selective labeling.
@@ 3.1 Conformal Labeling
Our previous section shows that existing methods cannot guarantee the quality of AI-assigned labels. To address this, we propose Conformal Labeling, a model-agnostic method to identify a subset where AI predictions can be provably trusted by controlling the FDR. Our method is composed of three primary steps: quantifying uncertainty, constructing conformal $p$-values, and thresholding.
@@
Our method builds on a key insight: we should select instances where the model exhibits high confidence in its predictions. To quantify the model confidence, we define an uncertainty (non-conformity) score $\mathcal{S}:\mathcal{X}\to\mathbb{R}$, where a higher value indicates greater model uncertainty. For example, we employ $\mathcal{S}(X)=1-\max_{y\in\mathcal{Y}}f_{y}(X)$ as our uncertainty score function (Hendrycks & Gimpel, 2016). This score is a valid measure of uncertainty, since prior works establish that misclassified samples generally receive lower probability scores (i.e., $\max_{y\in\mathcal{Y}}f_{y}(X)$) than correctly classified ones (Hendrycks & Gimpel, 2016; Tu et al., 2024).
@@
$H_{j}^{0}:\widehat{Y}_{n+j}\text{ is incorrect}\quad\text{v.s.}\quad H_{j}^{1}:\widehat{Y}_{n+j}\text{ is correct}.$
The correctness of AI-assigned labels depends on the task definition. For image labeling tasks, correctness is defined as $\widehat{Y}_{n+j}=Y_{n+j}$. For open-ended verifiable tasks, correctness corresponds to the generated output being a correct solution. For alignment tasks, correctness means that the generated output aligns with human preferences.
In the formulation, it is worth noting that rejecting the null hypothesis $H_{j}^{0}$ indicates that $(X_{n+j},\widehat{Y}_{n+j})$ should be included in the selection subset, as $\widehat{Y}_{n+j}$ is deemed to be correct. To construct the selection subset, we employ conformal p-value, which builds upon the conformal inference framework (Vovk et al., 1999, 2005). The underlying intuition is that: a test instance $X$ is likely to be misclassified if its uncertainty score $\mathcal{S}(X)$ is generally larger than the scores of instances that are known to be misclassified. Leveraging this idea, conformal $p$-value is computed through a rank-based comparison of $\mathcal{S}(X)$ against uncertainty scores of misclassified instances.
Formally, for the calibration dataset $\mathcal{D}_{cal}$, we identify the subset $\mathcal{D}_{cal}^{0}\subseteq\mathcal{D}_{cal}$ where instances are misclassified by the AI model. For simplicity, we denote $\mathcal{D}_{cal}^{0}=\{(X_{i},Y_{i})\}_{i=1}^{n_{0}}$, and thus $\widehat{Y}_{i}$ is incorrect for $i=1,\cdots,n_{0}$. We compute the uncertainty scores for the entire dataset $\{(X_{i},Y_{i})\}_{i=1}^{n+m}$: $\mathcal{S}_{i}=1-\max_{y\in\mathcal{Y}}f_{y}(X_{i})$. Then, the conformal $p$-value for the instance $X_{n+j}$ is computed by
$\widehat{p}_{j}=\frac{\sum_{i=1}^{n_{0}}\mathbf{1}\{\mathcal{S}_{i}<\mathcal{S}_{n+j}\}+U_{j}\left(1+\sum_{i=1}^{n_{0}}\mathbf{1}\{\mathcal{S}_{i}=\mathcal{S}_{n+j}\}\right)}{n_{0}+1},$ (3)
where $U_{j}\sim\mathrm{Uniform}[0,1]$ are i.i.d. uniform random variable to braek ties. Here, $\widehat{p}_{j}$ quantifies how extreme the uncertainty of $X_{n+j}$ is compared to the scores of mislabeled instances, with a small $\widehat{p}_{j}$ providing strong statistical evidence for correct prediction.
The conformal $p$-value defined in (3) has a distribution-free finite-sample guarantee, which is formalized in the following proposition. The conservatism property indicates that the conformal $p$-values for misclassified instances are biased to be high as they stochastically dominate the uniform distribution on $[0,1]$. This allows us to set a threshold to flag potential correct instances, while controlling the FDR.
Proposition 3.1.
(from Vovk et al. (2005)). Given that the calibration sample $\{(X_{i},Y_{i})\}_{i=1}^{n}$ together with the $j$-th test sample $(X_{n+j},Y_{n+j})$ are i.i.d. for $j\in\{1,\dots,m\}$, conformal $p$-value $\widehat{p}_{j}$ defined in (3) is conservative:
$\mathbb{P}\{\widehat{p}_{j}\leq\alpha\mid H_{j}^{0}\text{ is true}\}\leq\alpha,\quad\textnormal{for all }\alpha\in(0,1).$ (4)
Thresholding.
After obtaining conformal $p$-values for test instances, we can naively apply the Benjamini–Hochberg (BH) procedure (Benjamini & Hochberg, 1995) as in conformal selection (Jin & Candès, 2023) to select a subset $\mathcal{R}$ for AI labeling while controlling the FDR at level $\alpha$. Concretely, let $p_{(1)}\leq\ldots\leq p_{(m)}$ denote the ordered statistics of the $p$-values; the rejection set of the BH procedure applied to the conformal $p$-values is $\mathcal{R}=\{j:\widehat{p}_{j}\leq\widehat{p}_{(j^{*})}\}$, where
$j^{*}=\max\left\{j:\widehat{p}_{(j)}\leq\frac{\alpha j}{m}\right\},$ (5)
[…]
To mitigate this conservatism issue, we estimate $\pi_{0}$ from the calibration dataset, building upon the intuition that when the calibration data and test data are i.i.d., the model should achieve comparable accuracy on both datasets. In particular, we estimate $\pi_{0}$ with $\frac{1+n_{0}}{1+n}$, and subsequently apply the BH procedure at level $\frac{1+n}{1+n_{0}}\alpha$. We summarize the complete procedure of Conformal Labeling in Algorithm 1, which combines all three steps described above.
@@
0: Calibration set $\mathcal{D}_{\mathrm{cal}}=\{(X_{i},Y_{i})\}_{i=1}^{n}$, unlabeled test instances $\{X_{n+j}\}_{j=1}^{m}$, pre-trained model $f$, target FDR level $\alpha\in(0,1)$
0: A selected subset $\mathcal{R}$ with FDR control
1:// Identify mislabeled calibration data
2: Identify $\mathcal{D}^{0}_{\mathrm{cal}}\subseteq\mathcal{D}_{\mathrm{cal}}$ with size $n_{0}$
3:// Step 1: Compute uncertainty scores
4:for $i=1$ to $n+m$ do
5:  Compute uncertainty score $\mathcal{S}_{i}$
6:end for
7:// Step 2: Construct conformal $p$-values
8:for $j=1$ to $m$ do
9:  Construct conformal $p$-value $\widehat{p}_{j}$ according to Eq. (3)
10:end for
11:// Step 3: Thresholding
12: Compute $j^{*}=\max\left\{j:\widehat{p}_{(j)}\leq\frac{\alpha j(n+1)}{m(n_{0}+1)}\right\}$, where $\widehat{p}_{(j)}$ is the $j$-th smallest $p$-value.
13:Return $\mathcal{R}=\{j:\widehat{p}_{j}\leq\widehat{p}_{(j^{*})}\}$
In Theorem 3.2, we establish that Conformal Labeling controls the FDR below the target level $\alpha$ even when it applies the BH procedure at level $\frac{1+n}{1+n_{0}}\alpha$. This result provides a formal guarantee that the AI-assigned labels within the selected subset are trustworthy, in the sense that the expected proportion of incorrect labels is bounded by $\alpha$. Because the BH procedure is applied at an elevated level, Conformal Labeling selects more samples than naively applying the BH procedure at level $\alpha$. Moreover, Conformal Labeling’s upper bound for FDR is tight: it converges to $\alpha$ as the calibration set size increases and equals $\alpha$ in the asymptotic limit. The proof is provided in Appendix A.2.